MOA-II Obsevation Log

Return to index
Previous day Next day

DATE: 2006/08/01
OBSERVER: T. Sumi K.Ohnishi
CONDITION: fine, Overcast
COMMENT: Alart target MOA 2006 BLG116 (gb10-4) was observed every 10 min. After the second dark, chip 7 is unstable with noise, D1235-D1238 can not be used. We try ./@end and ./setup, after that chip 7 is stable.
Compressor gas pressure: MPa
Expected Observation Time: START 18:45 END 07:02


TARGETRUNTOTAL RUNTOTAL FIELDTOTAL EXPOSURE (sec)
GB3093-31273523313980
LMC-000
SMC-000
ALERT-000
GRB-000
TOTAL---3523313980 (03h53m00s)


GB FIELD STATISTICS

Field1Field2Field3Field4Field5Field6Field7Field8Field9Field10Field11Field12
888835888353277
Field13Field14Field15Field16Field17Field18Field19Field20Field21Field22Field23
76666666660


LMC FIELD STATISTICS

Field1Field2Field3Field4Field5Field6Field7Field8Field9Field10Field11Field12
000000000000
Field13Field14Field15Field16
0000


SMC FIELD STATISTICS

Field1Field2Field3Field4Field5Field6Field7Field8Field9Field10Field11Field12
000000000000


SPECIAL TARGET NAME

CategoryTargetNrun


CHECK SHEET

TimeStatusWeather
18:37observingfine, start the observation
19:30observingfine
20:30observingfine
21:30observingfine
22:30observingcloudy
23:30waitingstop due to cloud
00:30observingcloudy
01:30observingcloudy
02:30waitingOvercast
03:30waitingOvercast
04:30waitingOvercast
05:30waitingOvercast
06:30waitingOvercast
06:53waitingOvercast


CCD TEMPERATURE Previous day Next day




VACUUM Previous day Next day




TELESCOPE TEMPERATURE Previous day Next day




Return to index

Previous day Next day